OpenStudy (azureilai):

can someone explain how to use Pascal's triangle for cubing -1+((3i)^1/2)?

OpenStudy (amistre64):

pascals triangle has rows of numbers that represent the coefficient of the terms in a bimonial expansion

OpenStudy (amistre64):

the row for a ^3 is 1 3 3 1 then just define the ab parts to multiply to those

OpenStudy (amistre64):

1 3 3 1 pascals row -1 1 -1 1 (-1)^3,2,1,0 .... whatever the 3i 1/2s turn out to be from 0,1,2,3 ----------- multiply down and add across

OpenStudy (azureilai):

So would it be \[1(1^{3}+\sqrt{3i})+3(1^{2}\times \sqrt{3i}^{0})+3(1^{1} \times \sqrt{3i}^{2})+1(1^{0} \times \sqrt{3i}^{2}\] ?

OpenStudy (amistre64):

\[1(-1)^{3}(\sqrt{3i})^0~+3(-1)^{2}(\sqrt{3i})^{1}~+3(-1)^{1}(\sqrt{3i})^{2}+1(-1)^{0} (\sqrt{3i})^{3}\] \[-1+3\sqrt{3i}~-3(3i)+3i\sqrt{3i}\] yeah, something like that

OpenStudy (azureilai):

ok thanks

OpenStudy (amistre64):

youre welcome